OpenStudy (anonymous):

Consider two lines L1: 6x-8y=5 and L2: y=3x/4 -5. A(3/2,1/2) is a point on L1. (a) Show that L1//L2. [Done] (b) If L3 is a straight line passing through A and perpendicular to L1, find the equation of L3. (c) Find the point of intersection of L2 and L3.

hartnn (hartnn):

could you find slope of L1?

OpenStudy (anonymous):

3/4?

hartnn (hartnn):

yes.

hartnn (hartnn):

now, the slope of line perpendicular to line with slope m is -1/m so whats the slope of line L3 ?

OpenStudy (anonymous):

(3/4)*L3=-1 slope of L3=-4/3

hartnn (hartnn):

correct, now u have both slope and a point, can u find equation ? y-y1 = m(x-x1)

OpenStudy (anonymous):

y-(3/2)=(-4/3)*(x-1/2) ?

hartnn (hartnn):

yup, simplif it.

OpenStudy (anonymous):

8x+6y-13=0?

hartnn (hartnn):

yes, thats is correct.

OpenStudy (anonymous):

wait

hartnn (hartnn):

to find point of intersection, just solve the 2 equations of line simulatneously...

OpenStudy (anonymous):

I think I have some steps calculate wrongly.....the answer is 8x+6y-15=0

hartnn (hartnn):

let me check again...

hartnn (hartnn):

u interchanged x1 and y1

hartnn (hartnn):

y-(1/2)=(-4/3)*(x-3/2)

OpenStudy (anonymous):

Oh yes....!!! wait a moment, let me change it

OpenStudy (anonymous):

got it.... continue please.

hartnn (hartnn):

to find point of intersection, just solve the 2 equations of line simultaneously... can u ?

hartnn (hartnn):

y=3x/4 -5 and 8x+6y-15=0 solve simultaneously.

OpenStudy (anonymous):

please wait

OpenStudy (anonymous):

x=6/5 y=-41/10 why so strange...

hartnn (hartnn):

i got, 18/5 and -23/10 .....

OpenStudy (anonymous):

well.....wait..

OpenStudy (anonymous):

yes, I calculate wrongly. you are correct.

hartnn (hartnn):

good, any more doubts ?

OpenStudy (anonymous):

no more about this question. thank you very much. :)

hartnn (hartnn):

welcome ^_^